OpenStudy (anonymous):

Write an algebraic equation for the following. A number added to itself six additional times is four less than the number squared. A. x^7=x^2-4 B. 7x=4-x^2 C. 7x= x^2-4 D. 6x=x^2-4

OpenStudy (anonymous):

I think it's D because it's the only one with the number 6 in its equation

OpenStudy (anonymous):

yea that would be it cuz it is x+x+x+x+x+x=x*x-4=6x=x^2-4

OpenStudy (anonymous):

thanks(: